OpenStudy (anonymous):

I Need guidance. Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1 fourthx + 2 I don't know how to start this problem at all. I need help and guidance

OpenStudy (anonymous):

Anyone? :/

OpenStudy (texaschic101):

look at your equation : y = 1/4x + 2...can you tell me the slope of this equation ?

OpenStudy (anonymous):

\[Y=\frac{ 1}{ 4}x+2\]

OpenStudy (anonymous):

1/4 right?

OpenStudy (texaschic101):

correct..it is 1/4....and a parallel line will have the same slope. now we use y = mx + b we already know slope(m) = 1/4 (-4,-2)....x = -4 and y = -2 now we just need to find the y intercept, which is b now sub y = mx + b -2 = 1/4(-4) + b -2 = -1 + b -2 + 1 = b -1 = b so your parallel equation is : y = 1/4x - 1 any questions ?

OpenStudy (anonymous):

Ohhhh !! I Got it now Thank you sooo much.

OpenStudy (texaschic101):

do not get this confused with perpendicular lines...there a little different

OpenStudy (anonymous):

So all your doing is plugging in the numbers right?

OpenStudy (texaschic101):

yep...and solving for b, the y intercept

OpenStudy (anonymous):

Okayy Thanks

OpenStudy (texaschic101):

sure thing :)